OpenStudy (anonymous):

If the first term in an arithmetic series is 3, the last term is 136, and the sum is 1,390, what are the first 3 terms?

OpenStudy (anonymous):

3, 10, 17

OpenStudy (anonymous):

How did you find this out?

OpenStudy (anonymous):

Using Arithmetic Sum equation Sn=(n/2)(a_0+a_n) 1390=(n/2)(3+136) n/2=10 n=20 Arithmetic Sequences A_n= A_0 + d(n-1) 136= 3 + d(20-1) 133=19d d=7